OpenStudy (anonymous):

is x squared - x divisible by 3

OpenStudy (anonymous):

no. x^2-x=x(x-1)

OpenStudy (anonymous):

take for example x=8. U get 8*7=56. 56 is not divisible by 3.

OpenStudy (unklerhaukus):

\[\large3|(x^2-x)\iff3|[x(x-1)]\implies 3|x~\lor~3|(x-1)\]